A contact lens is made of plastic with an index of refraction of 1.32. The lens has an outer radius of curvature of +2.01 cm and an inner radius of curvature of +2.45 cm. What is the focal length of the lens? cm
Answer #1

The focal length of a lens is given by

                      1/f = (n-1)[1/R1 -1/R2]

                            =(1.32 -1)[1/2.01 -1/2.45]

                           = (0.32)(0.08935)

                           =0.028592

f= 1/0.028592

f=35 cm